Circular No. 8861

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 P1 (McNAUGHT) R. H. McNaught reports his discovery of a comet on 20-s CCD frames

obtained with the 0.5-m Uppsala Schmidt telescope at Siding Spring Observatory (discovery observation tabulated below), the comet showing a coma diameter of ∼ 40′′, extended to the southwest; McNaught adds that this is the most diffuse comet that he can remember observing, with little or no condensation and thus difficult to measure. Follow-up 60-s exposures by G. J. Garradd with the same telescope on Aug. 8.7 UT show some condensation, but the images are still very difficult to measure. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, J. Broughton (Reedy Creek, Qld., Australia, 0.51-m f/2.7 reflector) writes that his CCD frames taken on Aug. 9.7 show a 30′′-diameter coma and a 2′ tail in p.a. 350^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Aug. 7.72005  1^{h}59^{m}58.^{s}59  −31^{o}40′36′′.8  18.7

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-P23.

T = 2007 Mar. 26.439 TT  ω = 97.^{o}689
Ω = 190.879 }2000.0
q = 0.63318 AU  i = 122.963

POSSIBLE NOVA IN VULPECULA S. Nakano, Sumoto, Japan, reports the discovery of a possible nova

(mag 9.5) by Hiroshi Abe (Yatsuka-cho, Matsue, Shimane-ken, Japan) on several 30-s survey frames (limiting mag about 11.5) taken around Aug. 8.54 UT using a Canon EOS Kiss Digital Camera (+ 35-mm f/2.8 Pentax lens), the new object measured by Nakano from Abe’s JPEG image to be at α = 19^{h}54^{m}24^{s}.3, δ = +20^{o}52′47′′ (equinox 2000.0; estimated uncertainty ± 8′′); Abe notes that nothing was visible at this location on frames taken on July 23, 31, and Aug. 4. Nakano forwards the following position end figures for the new object from K. Kadota (Ageo, Saitama-ken, Japan, 0.25m reflector + unfiltered CCD, Aug. 8.682, mag 9.4): 24^{s}.64, 51′′.9; Kadota adds that a star of red mag 18.1 and blue mag 17.6 in the USNO-A2.0 catalogue has position end figures 24^{s}.66, 51′′.7. Visual magnitude estimates forwarded by E. Waagen, AAVSO: Aug. 9.006, 9.3 (S. Swierczynski, Dobczyce, Poland); 9.093, 8.7 (M. Komorous, London, ON, Canada); 9.188, 8.4 (R. King, Duluth, MN); 9.208, 8.7 (J. Bortle, Stormville, NY).

2007 August 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green